lizzie777:

The table contains some points on the graph of an exponential function. Capture.PNG" Determine the a and b values and formulate a model exponential equation to represent the data in the table. Based on the table, which value of a and b and which function represent the same relationship?

lizzie777:

1 attachment
lizzie777:

someone pls help me.....

lizzie777:

i dont know anything.........

AZ:

I'm assuming you also have answer choices? Regardless, exponential functions are in the form of \(y = ab^x\) we have to plug in the values given to us in the table to figure out what a and b are

AZ:

so the first point, we have x = 0 and y = 6 plug it in \( 6 = ab^{0}\)

AZ:

do you know what anything to the power of 0 is?

lizzie777:

i think so

AZ:

anything to the power of 0 is 1 so what is b^ 0 =

lizzie777:

B^1?

AZ:

what is b^0 = I just told you that anything to the power of 0 equals 1 so 10^0 = 1 a^0 = 1 what is b^0 =

AZ:

it's not a trick question

lizzie777:

=1

AZ:

Good! \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so the first point, we have x = 0 and y = 6 plug it in \( 6 = ab^{0}\) \(\color{#0cbb34}{\text{End of Quote}}\)

AZ:

So now we have 6 = a * 1 what is a equal to?

lizzie777:

would it be 6 bc its multiplying by 1?

AZ:

yes! so now that we know a, let's solve for b so we have \(\Large y = 6b^x\)

AZ:

Let's plug in the next point (1, 9) x = 1 and y = 9 \(\Large 9 = 6 \times (\text{b}^1)\) what is b^1 =

AZ:

anything to the power of 1 is the same thing so like 13^1 = 13

lizzie777:

so its 1

AZ:

no...

AZ:

what is b to the power of 1 b^1 = anything to the power of 1 is that same thing ex. d^1 = d

lizzie777:

oh ...B

AZ:

yes so now we have 9 = 6 * b what is b = ?

lizzie777:

......6

AZ:

what number times 6 will give you 9

AZ:

b times six equals 9 b * 6 = 9 we want to move six to the other side divide it on both sides then

lizzie777:

well 6 divied by 9 is 0.66

AZ:

no

lizzie777:

or 1.5

AZ:

look let me change the letter because b is looking a whole lot like 6 let's make it F F * 6 = 9

AZ:

if we're dividing 6 on both sides

AZ:

do you do 6/ 9 or 9/ 6 ???????????????

AZ:

we want to get the letter all by itself so we divide by 6 on both sides

lizzie777:

well 9 divied by 6 is 1.5

AZ:

there you go

AZ:

so that's the value of b

AZ:

now what's the final correct answer? we said a was 6 and b was 1.5 \(\Large y = ab^x\)

lizzie777:

y=6a(1.5b)\[y=6a1.5b ^{x}\]

AZ:

no, you get rid of the a and the b and put the numbers instead do you see your answer choices having an `a` and `b` ????????????

lizzie777:

6(1.5)x

AZ:

there you go (I know you copy pasted but otherwise it should be (1.5)^x )

lizzie777:

so the answer choice is B

AZ:

yes

lizzie777:

thank you!

AZ:

you're welcome!